1. A compound of carbon, Hydrogen, Nitrogen and Oxygen was analyzed. Its chemical composition was determined as follow: 49.5% C, 5.2% H, 28.8% N and 16.5% O.

(a) Determine the empirical formula of the compound.
(b) If 0.050 mol of the compound referred to in (a) weight 9.71 g. Find the molecular mass of the compound.
(c) Workout the molecular formula of the compound.
1
Expert's answer
2020-08-29T14:02:15-0400

a) Considering the given percentages, 100 g of the compound contains:

49.5 g or 49.5 / 12.01 = 4.12 mol of Carbon

5.2 g or 5.2 / 1.01 = 5.15 mol of Hydrogen

28.8 g or 28.8 / 14.01 = 2.06 mol of Nitrogen

16.5 g or 16.5 / 16.00 = 1.03 mol of Oxygen

The C : H : N : O mole ratio is 4.12 : 5.15 : 2.06 : 1.03 = 4 : 5 : 2 : 1

Therefore, the empirical formula is C4H5N2O

b) M = 9.71g / 0.050mol = 194.2 g/mol

c) The molecular mass of one empirical formula unit is 12.01*4 + 1.01*5 + 14.01*2 + 16 = 97.1 g/mol.

The molecular mass of the given compound is twice the molecular mass of one empirical formula unit: 194.2 / 97.1 = 2. Therefore, subscripts of the empirical formula should be multiplied by 2. The formula needed is C8H10N4O2.